### v3

# Large Gravitational Wave Phase Shifts from Strong 3-body Interactions in Dense Stellar Clusters

Abstract
The phase evolution of gravitational waves (GWs) can be modulated by the astrophysical environment surrounding the source, which provides a probe for the origin of individual binary black holes (BBHs) using GWs alone. We here study the evolving phase of the GW waveform derived from a large set of simulations of BBH mergers forming in dense stellar clusters through binary-single interactions. We uncover that a well-defined fraction of the assembled eccentric GW sources will have a notable GW phase shift induced by the remaining third object. The magnitude of the GW phase shift often exceeds conservative analytical estimates due to strong 3-body interactions, which occasionally results in GW sources with clearly shifted and perturbed GW waveforms. This opens up promising opportunities for current and future GW detectors, as observing such a phase shift can identify the formation environment of a BBH, as well as help to characterise the local properties of its surrounding environment.
preprint: APS/123-QED
I Introduction
Gravitational wave (GW) templates for binary black hole (BBH) mergers are constructed with the masses, spins, and orbital eccentricities of the progenitor systems. The different proposed formation channels of BBH produce distinct features in these three parameters, indicative of the BBH formation mechanism [e.g. 103, 81, 28, 67, 68, 72, 74, 63, 65, 61, 42]. While powerful, this method only allows for population-level statements and cannot on its own directly show how a single BBH formed.
A novel approach that is gaining increasing attention is to identify direct modulations of the GW signals of binaries due to the astrophysical environment in which they form, which, if observable, yield a means to probe the environment and formation mechanism of the BBH itself on a single-event basis [e.g. 98]. These modulations include general relativistic propagation effects (e.g., GW source acceleration [e.g. 102, 48, 36, 57, 19, 56, 100, 84, 20, 88, 80, 101, 39, 98, 71, 30], GW lensing and gravitational redshift [e.g. 27, 53]), astrophysical environmental effects (e.g., gas dynamical friction [e.g. 11, 105] and tidal forces [71]), and effects beyond classical general relativity (GR) [e.g. 17, 18]. As no BBH merges in a completely empty Universe, such effects should leave imprints in every observed GW signal; however, in most cases the modulations are too small to be resolved. This has led to the key question whether any of the considered BBH merger formation channels such as stellar clusters [54, 40, 10, 85, 9, 62, 55, 59, 60, 60, 7, 51, 61, 69, 65, 70, 90, 90, 91, 93], isolated binary stars [21, 22, 23, 14, 13, 76, 49, 58, 75, 77, 86, 87, 37], active galactic nuclei (AGN) discs [12, 79, 47, 82, 64, 94, 25], or galactic nuclei [50, 32, 95, 6, 78, 31, 29, 92, 41, 43, 8, 89], naturally produce a notable fraction of mergers with measurable GW modulations able to inform about the environment of the merger.
Figure 1: Illustration of 3-body interaction resulting in a BBH merger with an observable GW Phase Shift. Top left: A stellar cluster with highlighted BH interactions, each of which is able to produce BBH mergers. Top right: Zoom-in on a binary-single interaction resulting in a BBH merger with the third object still bound (3-body merger). Turquoise- and pink lines show the trajectory of the merging BHs (true path), where the orange lines illustrate the path the BBH would have taken without the third object (reference path). The trajectory of the third BH is depicted in red. The white lines show lines-of-sight for an observer located in the lower right corner, where the green lines illustrate the spacial distance between the true- and the reference paths along the sight-lines, respectively. Bottom: GW strain as a function of time. The turquoise curve shows the GW signal for the observed BBH inspiral (true path), where the orange shows what the isolated BBH merger signal would look like (reference path). The two signals are shifted by the light crossing time between the true and the reference BBH paths, i.e. the time it takes the GWs to travel along the green lines. This gives rise to a unique observable GW phase shift that can be directly mapped to the BBH formation and environment.
In this Letter we perform post-Newtonian ($\cal PN$}) N-body simulations (see Methods), to show that BBHs formed through binary-single interactions occurring in dense stellar clusters naturally lead to a well-defined population of GW sources with a measurable GW phase shift caused by the presence of the third object. This is illustrated in Fig. 1, which shows an illustration of different dynamical interactions in a cluster environment (upper left), a BH binary-single interaction producing a GW-driven merger (upper right), and the corresponding GW phase-shifted signal caused by the presence of the third-object (lower panel). Several dynamical pathways can lead to BBH mergers in clusters [e.g. 70], but the pathway relevant for clearly producing phase-shifted sources observable by ground based detectors is characterized by two of the three interacting BHs inspiralling and merge while the third is still bound (Fig. 1 upper right and Fig. 2), also referred to as a 3-body merger (see Methods). Such 3-body mergers constitute $\sim 10\%$ of the BBH mergers from globular clusters (GCs) [69] and dominate the fraction of eccentric sources, which implies that GW phase-shifted sources will form in notable numbers with observable prospects for both current ground-based GW detectors (LIGO/Virgo/KAGRA) [1, 3, 4, 2] and future (Einstein Telescope (ET) [46], Cosmic Explorer (CE) [24]), as well as space-borne detectors (DECIGO/TianQin/Taiji [38, 45, 33, 44] (deci-Hertz) and LISA [5] (milli-Hertz)).
Figure 2: Distribution of maximum GW phase shifts from 3-body BBH mergers. Results from $\cal PN$} simulations between a BBH and a single incoming BH that concludes with two of the three BHs merging while the third BH is still bound (see Fig. 1 and Fig. 3). Each dot shows the maximum observable GW phase shift (y-axis) for each of these mergers, as a function of the corresponding GW peak frequency (bottom x-axis) or merger time (top x-axis), derived using the methods outlined in Sec. II. Each colour refers to a different SMA of the initial BBH before interaction: $\sim 1$ AU (orange), $\sim 0.1$ AU (red), and $\sim 0.01$ AU (blue). The dashed coloured lines illustrate the asymptotic limit $\Delta{\phi}\propto f^{-13/3}$, where the horizontal dotted line indicates the analytical maximum value for $\Delta{\phi}$ assuming the distance between the BBH and the single BH, $R$, is similar to the initial SMA, $a$ (see Methods). The three highlighted examples, (A,B,C), are shown and studied further in Fig. 3.
II Gravitational Wave Phase Shifts
The GW phase shift in 3-body mergers arises from the BBH being bound to the single BH while merging, and therefore moving on a curved orbit with a time-varying velocity along the line-of-sight (LOS) relative to the observer. This creates a time-dependent Doppler shift of the GW waveform, which translates to a GW phase shift that can be mapped to the dynamics of the 3-body system [71], analogous to binary pulsar timings [e.g. 48, 71].
In the limit where the acceleration $\bf a$} of the BBH center-of-mass (COM) near merger from the presence of the third body can be assumed constant, the maximum displacement in time between a GW signal sent from the accelerated BBH path and a non-accelerated path as seen by a distant non-cosmological observer is given by [e.g. 30],
$\tau(t)=\frac{1}{2c}|{\bf a}|t^{2}=\frac{1}{2}\frac{Gm}{c}\frac{t^{2}}{R^{2}}.$ (1)
Here we assume all three BHs have equal mass $m$, $c$ denotes the speed-of-light, and $R$ refers to the distance between the BBH COM and the single BH at the time of merger (see Methods for solution to the unequal mass case). This $\tau$ corresponds to the time it takes a light pulse to travel along one of the green lines shown in Fig. 1; a time that often is referred to as the Rømer Delay. From $\tau$, the GW phase shift $\phi$, which is the difference in angular cycles of the GW signal in units of radians, can now be expressed as [71] {align} Δϕ(t) ≈2πτ(t)T(t) ≈22G3/2c m3/2R2 ×t2a(t)3/2  , where $T(t)$ and $a(t)$ denote the BBH orbital time and the BBH semi-major axis (SMA), respectively, at time $t$. Using the analytical framework of [52], hereafter referred to as Peters64, for relating the time $t$ to the BBH’s eccentricity, $e$, and SMA, $a$, as well as the relation for $a(e)$, one now finds (see Methods), {align} Δϕ(e) ≈144852g(1)13/2 c9G9/2 ×1R2r013/2m9/2
×e^78/19(1-e^2)^1/2g(e)^13/2, where $g(e)=\left(1+121e^{2}/304\right)^{870/2299}$, and $r_{0}$ is the peri-center distance of the merging BBH at the time of formation during the interaction. This $r_{0}$ maps to a corresponding GW peak frequency as [e.g. 69]
$f_{0}\approx\frac{1}{\pi}\sqrt{{2Gm}/{r_{0}^{3}}},$ (2)
[…]
$f_{m}\approx{f_{0}}\left(\frac{2e_{m}^{12/19}}{1+e_{m}}\frac{g(e_{m})}{g(1)}\right)^{-3/2}.$ (3)
[…]
III Results
Figure 3: Examples of 3-body BBH mergers resulting in significant GW phase shifts. The shown interactions involve BHs with equal mass $m=20M_{\odot}$, and initial SMA $a=0.1$ AU for case (A) and (B), where case (C) has $a=0.01$ AU. The filled-circles in the left figures indicate the initial positions of the BHs, where the filled-circles in the right figures show the end positions at merger. The orange lines illustrate the trajectory the merging BBHs would take if the third BH was not there (see Fig. 1). Common for the cases that give rise to significant GW phase shifts is that the chaotic nature of the 3-body problem brings the BBH close to the remaining bound single BH near merger. This is clearly seen in these examples. GW phase shifts and corresponding peak frequencies for (A,B,C) are shown in Fig. 2.
We now consider results from controlled $\cal PN$} binary-single interactions between BBHs with initial SMA $a=0.01,0.1,1~\mathrm{AU}$}, respectively, and an incoming BH, all with equal mass $m=20~M_{\odot}$. These values are characteristic for BHs interacting in dense stellar systems, as these map to stellar clusters with escape velocities, $v_{e}$, of $\sim 30\ \mathrm{km\,s}^{-1}$ (globular cluster), $\sim 90\ \mathrm{km\,s}^{-1}$ (dense stellar cluster), $\sim 300\ \mathrm{km\,s}^{-1}$ (nuclear star cluster), for $a=1,0.1,0.01~\mathrm{AU}$, respectively [e.g. 6]. For every 3-body merger that forms, we measure numerically $r_{0}$ and $R$, from which we estimate the maximum GW phase shift and corresponding GW peak frequency, as outlined in Section II.
Fig. 2 shows our numerically derived distributions of $f_{m}$ and $\Delta{\phi}(e_{m})$. As seen, the chaotic nature of the binary-single problem leads to GW sources with notable GW phase shifts, especially in the range between 1101-10 Hz, accessible to third-generation (3G) ground-based GW observatories such as ET and CE. For instance, $\sim 20\%$ of the population with $a_{0}=0.1$ AU (red) has a maximum phase shift above 0.01 radians. This result is non-trivial, as an analytical estimate for the maximum GW phase shift results in a near universal upper limit of $\Delta{\phi}_{H}(e)\sim 10^{-2}\times(m[M_{\odot}]/R[\mathrm{AU}])^{1/7}$ radians [71], which is derived using Eq. 1 with $r_{0}$ set to its maximum possible value, corresponding to where the merging BBH starts exactly at its Hill sphere relative to the single BH at distance $R$ (see Methods).
To understand what kind of 3-body evolutions are able to bypass the analytical limit of $\sim 10^{-2}$ radians, we now consider the three distinct examples shown in Fig. 3. In (A), we show a highly chaotic and long-lived interaction, that concludes with an endstate charaterized by the BBH passing the bound single near closest approach on a clearly curved orbit. In (B), we show an example where the BBH merges right after turning around to directly move towards the bound single. This creates a large offset between the reference- and the true trajectory. Note also how the single BH scatters off the BBH near the 3-body COM $(0,0)$ as it inspirals. This creates large extra perturbations to the GW signal. In (C), we show a case where the merging BBH is heading directly towards the bound single BH. Here there is no significant curvature, but the acceleration along its direction of motion towards the single BH creates the GW phase shift. Note that the orbital precession due to the 1- and 2-$\cal PN$} terms is included in this figure for a more accurate illustration of the dynamics.
As follows from Eq. 1, the only important parameter that can freely vary during interactions, and also lead to a large GW phase shift at a given $f$, is the distance $R$. The GW sources we see forming with the largest GW phase shift are therefore all characterized by having the remaining single BH much closer at merger than expected from the initial binary separation. As illustrated, this all follows naturally from the chaotic nature of the 3-body problem, which further opens up for new directions in performing, classifying and identifying GW phase-shifted sources in $\cal PN$} few-body scatterings.
GW phase shifts being a natural outcome when BBHs merge through few-body interactions in stellar clusters opens up tremendous possibilities for probing the origin and assembly mechanisms of individual GW sources. In particular, many of these systems form at GW frequencies of a few hertz, placing them near the lower end of the sensitivity band of future observatories such as ET and CE. GW phase shifts from Rømer delay have been quantified to be among the loudest type of environmental effect [106, 83, 107] and, given their large amplitude found in this work and the sensitivity of next-generation detectors, they are expected to be present in a significant portion of GW signals if these dynamical environments contribute to the merger rate.
Acknowledgements.
The authors are grateful to Lorenzo Speri, Fabio Antonini and Daniel J. D’Orazio for useful discussions. K.H, L.Z., P.S., J.T, and J.S. are supported by the Villum Fonden grant No. 29466, and by the ERC Starting Grant no. 101043143 – BlackHoleMergs led by J. Samsing. M.Z. gratefully acknowledges funding from the Brinson Foundation in support of astrophysics research at the Adler Planetarium. A.A.T. acknowledges support from the Horizon Europe research and innovation programs under the Marie Skłodowska-Curie grant agreement no. 101103134.